### v1

@@ Abstract
A new generation of neutrino observatories, including RNO-G and RET, will search for PeV-EeV neutrinos interacting in the ice by detecting radio pulses. Extended air showers propagating into the ice will form an important background and could be a valuable calibration signal. We present results from a Monte-Carlo simulation framework developed to fully simulate radio emission from cosmic-ray particle cascades as observed by in-ice radio detectors in the polar regions. The framework involves a modified version of CoREAS (a module of CORSIKA 7) to simulate in-air radio emission and a GEANT4-based framework for simulating in-ice radio emission from cosmic-ray showers as observed by in-ice antennas. The particles that reach the surface of the polar ice sheet at the end of the CORSIKA 7 simulation are injected into the GEANT4-based shower simulation code that takes the particles and propagates them further into the ice sheet, using an exponential density profile for the ice. The framework takes into account curved ray paths caused by the exponential refractive index profiles of air and ice. We present the framework and discuss some key features of the radio signal and radio shower footprint for in-ice observers.
@@ I Introduction
Radio signals have an attenuation length 10 times longer than optical light in polar ice, making them a favourable means to monitor very large interaction volumes, and thus enable UHE neutrino detection in the $>$ PeV energy range [5, 6, 7, 8, 9]. Current in-ice radio neutrino detection experiments with typical peak sensitivities in the higher PeV to EeV energy range consist of two classes. The first class of experiments includes projects like ARA [10], ARIANNA [11], and RNO-G [12]. These experiments have deployed radio antennas inside the Antarctic and Greenlandic ice sheets and aim to detect coherent radio emission, known as Askaryan emission, from particle cascades induced by neutrino interactions in ice [13, 14, 15].
The second class of experiments aims to detect UHE neutrino particle cascades using radar echoes. The in-ice particle cascades are illuminated with radio waves and the reflected signal is recorded, which could potentially cover the PeV - EeV energy gap. The T-576 experiment at SLAC recently demonstrated that the radar echo method works for particle cascades in High-Density PolyEthylene (HDPE), a material having very similar properties to polar ice [16]. Currently, the Radar Echo Telescope for Cosmic Rays (RET-CR) has been deployed at Summit Station, Greenland, as a pathfinder experiment to test the radar echo method in nature [17].
@@
Radio emission from UHE Cosmic Ray (UHECR) particle cascades acts as an important background for in-ice radio neutrino detectors. UHECRs have a much higher interaction probability and arrive with a considerably higher flux than UHE neutrinos, while the radio signals emitted from particle cascades induced by cosmic rays in air and ice share many similarities with those emitted from in-ice neutrino-induced particle cascades. Furthermore, they can serve as an important proof-of-concept and could be an interesting calibration source for the in-ice radio detectors. It is therefore critical to understand and properly simulate the radio emission from cosmic-ray particle cascades as seen by in-ice detectors. In this work we present FAERIE, the Framework for the simulation of Air shower Emission of Radio for in-Ice Experiments. It is the first complete Monte-Carlo cosmic-ray radio emission simulation framework for in-ice detectors, including both the propagation of the particle cascade in air and in ice, and discuss the expected emission and its main features.
@@ II Framework overview
To fully simulate the radio emission of cosmic-ray particle cascades for in-ice radio antennas, we need to consider: a) the emission created during the development of the in-air particle shower and how this propagates into ice, and b) the emission created during the development of the in-ice particle cascade, initiated when the particle shower itself propagates into ice. Earlier studies investigating the propagation of cosmic-ray particle cascades into ice and other media can be found in [22, 23, 24, 25, 26, 27, 28, 29, 30]. Since polar ice can reach up to altitudes of $\mathcal{O}$(3 km), where UHECR air showers still have a very energy-dense core, the in-ice radio emission can lead to a significant contribution to the total radio emission of the shower as observed by in-ice antennas.
@@
FAERIE uses the air shower simulation program CORSIKA 7.7500 to simulate the in-air particle cascade, which calculates the radio emission using the CoREAS code [31, 32]. CoREAS applies the endpoint formalism to simulate the electric field at a given antenna position in air. This is done by calculating the emission of every single charge in the particle cascade during the CORSIKA simulation, and uses straight-line propagation of rays [33, 34]. For FAERIE, we modified the code so it can handle antenna positions in ice using full ray tracing. The ray tracing has been discussed in detail in  [35, 36].
@@
For the simulation of the in-ice particle cascade, FAERIE uses a code based on the GEANT4 simulation toolkit [37]. GEANT4 allows for full Monte-Carlo simulation of particle creation and propagation through any given medium. Using GEANT4 we propagate the core of the in-air particle shower, i.e. all particles within $1$ m of the shower axis, into an ice volume consisting of multiple horizontal 1-cm-thick layers of pure ice. As illustrated by Figures 16 and 17 in Appendix A, most of the energy deposited in the ice is well contained within a radius of $1$ m. This is also shown in Figures 5 and 6 in [29]. Furthermore, Figure 16 in [29] shows that we do not expect any significant changes in the total fluence of the in-ice radio emission around the Cherenkov cone, when the radius of the particle footprint being propagated into the ice is increased beyond $\sim 10$ cm. Each layer of the ice has a constant density determined by the depth of the layer, following a typical polar ice density profile given by
@@
[…]
@@ III Ray tracing
For the results presented in this work, the density of the atmosphere is described by a five-layer model for the South Pole fitted to a database used by the National Center for Environmental Predictions (NCEP) Global Forecast System (GFS) in weather forecasting. Based on this fit, a five-layer model of the refractive index is created. Both the five-layer density model as well as the corresponding five-layer refractive index model are then used for the CORSIKA/CoREAS simulation [35]. The details of the models are given in Appendix B. In the following, the refractive index model for ice is set to
@@
corresponding to the model used by the Askaryan Radio Array (ARA) at the South Pole outlined in [40]. More detailed information on the analytical ray tracer can be found in [41, 42].
The ray-bending effect is less pronounced for air than for ice since the air refractive index value remains close to unity throughout. The refractive index profile is however still relevant, as the refractivity $R=n-1$ starts around $R=0$ at the top of the atmosphere and increases to approximately $R=3\times 10^{-4}$ at sea level, which can lead to a so-called refractive displacement for inclined air showers [43]. The polar ice refractive index changes from $n=1.35$ at the surface to $n=1.78$ within the first $100-200~{}$m. Therefore, significant ray bending occurs near the ice surface due to a sharp change in the refractive index values. Moreover, for ice-to-ice ray propagation, in general two solutions are found for each emitter-receiver pair. The first solution we will call the direct ray, which out of the two solutions has the shortest ray path between emitter and receiver. The second solution we will call the indirect ray, which can either be a refracted ray with a ray path longer than that of the direct ray, or a ray reflecting on the ice-air boundary. All three types of rays are shown in Figure 1.
@@ III.1 Interpolation of ray tracing parameters
A typical execution of the analytic ray tracing functions involving air and ice takes around $0.05-0.1$ ms. This is relatively fast but not fast enough for simulating a complete cosmic ray shower. A typical cosmic-ray shower will consist of $\mathcal{O}(10^{9})$ particles at EeV energies. In this case, propagating rays from each shower particle at each step to each antenna could take weeks or months, depending on the shower energy and the number of in-ice antennas. Therefore, it is not feasible to use the analytic ray tracing functions directly, and we have to move towards interpolation tables. Interpolation of ray tracing parameters from pre-made tables makes the ray tracing process significantly faster.
@@
To reduce memory usage during the simulation, antennas at the same depth in the ice share a single interpolation table for the in-air ray tracing. Since the index of refraction profile only depends on depth, the ray tracing applied is cylindrically symmetric. A grid covering a sufficiently large interval of horizontal distances can therefore be used by multiple antennas located at the same depth in the ice. For the in-air ray tracing the grid depth is varied from the ice layer altitude, generally around 3 km, up to 100100100~{}100km. The horizontal distance for the grid is varied by varying the ray launch angle ($\theta$) from $89.9^{\circ}$ (almost horizontal) to $0^{\circ}$ (vertically down). Small variations in $\theta$ can cause almost exponential changes in the horizontal distance.
The in-ice ray tracing interpolation tables span intervals of 20 m on both axes, which covers the typical dimension of in-ice cascade. Since in ice the refractive index changes much more than it does in air, the in-ice grids are denser than the in-air grids. For in-ice ray tracing, one interpolation table is made for every single antenna, irrespective of the antenna depths in the ice. A visualisation of an interpolation table used for the in-ice ray tracing is given in Figure 2.
@@
Figure 2: A visualisation of an interpolation table used for the in-ice ray tracing, for a given receiver position. Interpolation tables for the in-air ray tracing follow a similar structure, but instead cover a much larger area in air. A single in-air table is used for multiple receivers at the same depth in the ice.
@@ III.2 Modification of the endpoint formalism
Left: Visualisation of several rays traced from the receiver to a line of emitters, including an air-ice boundary at $z=0$. Direct rays are shown in blue. Reflected rays are only considered when total internal reflection occurs and are shown in red. Transmitted rays are shown in green. Right: The boost factor estimations along with the numerically computed value of $\frac{dt}{dt^{\prime}}$ as a function of depth. The diamonds correspond to the calculations using $\hat{r}$ pointing directly from emission point to receiver point. The stars correspond to the values of the boost factor when evaluating $n$ at the emission point and using the launch direction for $\hat{r}$. The dots correspond to the numerically calculated values of the boost factor. The different regions are associated with direct (blue), reflected (red) and transmitted rays (green).
Figure 3:
The end-point formalism calculates the electric field observed at a position $\vec{x}$ emitted by a charged particle taking a propagation step (track segment) in the simulation by using the formula
@@
where the plus sign is applied using the start point of the step and the minus sign using the end point of the step [33]. The combination of both then gives the net contribution of the step. Here, $\Delta t$ is the sampling time interval of the observer, $q$ the charge of the particle, $\hat{r}$ the direction from the start/end-point towards the antenna, $\vec{\beta}^{*}$ the velocity of the charge during the step, $R$ the distance between the start/end-point and the antenna and $n$ the index of refraction of the medium at the respective emission point. The variable $t$ refers to the arrival time of the emission at the observer, and relates to the emission time $t^{\prime}$ via $t=t^{\prime}+nR/c$. The emission time $t^{\prime}$ is often referred to as the retarded time.
The so-called boost factor $B=1-n\vec{\beta}^{*}\cdot\hat{r}$ found in Equation 3 arises from classical theory and can more generally be described as $B=\frac{dt}{dt^{\prime}}$. The implementation as used in the endpoint formalism considers signals traveling along straight lines, i.e. $\hat{r}$ points directly from emission point to receiver point. This approach is however not valid when working in media with a non-constant index of refraction where formalisms such as ray tracing need to be used to correctly describe signal propagation.
Using a ray tracer calculating the derivative $\frac{dt}{dt^{\prime}}$ numerically, we found that the relation $B=1-n\vec{\beta}^{*}\cdot\hat{r}$ still holds provided that $n$ is evaluated at the emission point and $\hat{r}$ is interpreted as the launching direction of the ray [44]. Figure 3 shows a visualisation of the ray tracing together with the comparison of different estimators of the boost factor $B$ to the numerically computed value of $\frac{dt}{dt^{\prime}}$. Rays were traced from a receiver in the ice to a line of emitting points, including an air-ice boundary at $z=0$. Both the in-air and the in-ice medium have an associated exponential index of refraction profile. Evaluating the boost factor using a straight line calculation clearly diverges from the numerically calculated value, while evaluating $n$ at the emission point and using the launch direction for $\hat{r}$ agrees well for all direct, reflected and refracted rays.
@@
Finally, a rotation of the electric field is applied in the plane of the ray. This rotates the electric field such that its component in the plane of the ray is perpendicular to the receiver direction, instead of the launching direction. A visualisation of the endpoint formalism including ray tracing is shown in Figure 4.
Figure 4: A visualisation of the variables used in the end-point formula including ray tracing (Equation 3), in the specific case where the start point, end point and receiver coincide in the same plane. The circles represent the start and end point of the emitter step, the square represents the receiver. All vectors shown in the illustration lie in the same plane.

### v4

@@ Abstract
A new generation of neutrino observatories will search for PeV-EeV neutrinos interacting in the ice by detecting radio pulses. Extended air showers propagating into the ice will form an important background and could be a valuable calibration signal. We present results from a Monte-Carlo simulation framework developed to fully simulate radio emission from cosmic-ray particle cascades as observed by in-ice radio detectors in the polar regions. The framework involves a modified version of CoREAS (a module of CORSIKA 7) to simulate in-air radio emission and a GEANT4-based framework for simulating in-ice radio emission from cosmic-ray showers as observed by in-ice antennas. The particles that reach the surface of the polar ice sheet at the end of the CORSIKA 7 simulation are injected into the GEANT4-based shower simulation code that takes the particles and propagates them further into the ice sheet, using an exponential density profile for the ice. The framework takes into account curved ray paths caused by the exponential refractive index profiles of air and ice. We present the framework and discuss some key features of the radio signal and radio shower footprint for in-ice observers.
@@ I Introduction
Radio signals have an attenuation length 10 times longer than optical light in polar ice, making them a favorable means to monitor very large interaction volumes, and thus enable ultra high energy (UHE) neutrino detection in the $>$ PeV energy range [5, 6, 7, 8, 9]. Current in-ice radio neutrino detection experiments with typical peak sensitivities in the higher PeV to EeV energy range consist of two classes. The first class of experiments includes projects like ARA [10], ARIANNA [11], and RNO-G [12]. These experiments have deployed radio antennas inside the Antarctic and Greenlandic ice sheets and aim to detect coherent radio emission, known as Askaryan emission, from particle cascades induced by neutrino interactions in ice [13, 14, 15].
The second class of experiments aims to detect UHE neutrino particle cascades using radar echoes. The in-ice particle cascades are illuminated with radio waves, and the reflected signal is recorded, which could potentially cover the PeV-EeV energy gap. The T-576 experiment at SLAC recently demonstrated that the radar echo method works for particle cascades in High-Density PolyEthylene, a material having very similar properties to polar ice [16]. Currently, the Radar Echo Telescope for Cosmic Rays has been deployed at Summit Station, Greenland, as a pathfinder experiment to test the radar echo method in nature [17].
@@
Radio emission from UHE Cosmic Ray (UHECR) particle cascades acts as an important background for in-ice radio neutrino detectors, as UHECRs have a much higher interaction probability and arrive with a considerably higher flux than UHE neutrinos. Furthermore, they can serve as an important proof of concept and could be an interesting calibration source. It is therefore critical to understand and properly simulate the radio emission from cosmic-ray particle cascades as seen by in-ice detectors. To fully understand the properties of UHECR radio emission, not only the emission during the development of the particle cascade in air has to be taken into account. Especially for detectors deployed at high altitudes, such as at the South Pole or near the summit of the Greenland ice sheet, the emission created during the propagation of the particle cascade through the ice should be considered as well, since at these altitudes the particles in the shower cascade still carry a significant fraction of the energy of the primary particle.
In a previous work, we have presented simulation results of the propagation of UHECR air showers in ice [22]. We used the CORSIKA Monte Carlo code [23] to simulate air showers, and developed a module based on the GEANT4 simulation toolkit [24] to propagate the air shower particle footprints through a high-altitude ice layer. We showed that the particle showers contain a very energy-dense core, forming the main component determining the development of the in-ice part of the cascades. We found that the lateral charge distribution of the in-ice cascades has a comparable typical width to that of neutrino-induced particle cascades, indicating that the radio emission from both types of events shares many similarities. We implemented the endpoint formalism [25] within the GEANT4 module to get a first estimation of the Askaryan emission from the in-ice particle cascades, ignoring detailed radio signal propagation in the ice. We concluded that the energy-dense core containing the particles of the cascade up to a radius of $1$ m dominates the radio emission, leading to similar signals compared to neutrino-induced particle cascades.
Following up on the previous work summarized above we now present FAERIE, the Framework for the simulation of Air shower Emission of Radio for in-Ice Experiments. It is the first complete Monte-Carlo cosmic-ray radio emission simulation framework for in-ice detectors. It includes both the propagation of the particle cascade in air and in ice, and applies the endpoint formalism to calculate the radio emission from both components using ray tracing to account for radio propagation through nonuniform media.
@@ II Framework overview
To fully simulate the radio emission of cosmic-ray particle cascades for in-ice radio antennas, we need to consider: a) the emission created during the development of the in-air particle shower and how this propagates into ice, and b) the emission created during the development of the in-ice particle cascade, initiated when the particle shower itself propagates into ice. Earlier studies investigating the propagation of cosmic-ray particle cascades into ice and other media can be found in [14, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 22, 36]. Polar ice can reach up to altitudes of $\sim 3$ km, which corresponds to a vertical atmospheric depth of $\sim 730$ g/cm2 at the South Pole, where UHECR air showers still have a very energy-dense core. As such, the in-ice radio emission can lead to a significant contribution to the total radio emission of the shower as observed by in-ice antennas.
@@
FAERIE uses the air shower simulation program CORSIKA 7.7500 to simulate the in-air particle cascade, which calculates the radio emission using the CoREAS code [23, 37]. CoREAS applies the endpoint formalism to simulate the electric field at a given antenna position in air. This is done by calculating the emission of every single charge in the particle cascade during the CORSIKA simulation, and using straight-line propagation of rays [25, 38]. For FAERIE, we modified the code so it can handle antenna positions in ice using full ray tracing. The ray tracing has been discussed in detail in  [39, 40].
@@
For the simulation of the in-ice particle cascade, FAERIE uses a code based on the GEANT4 simulation toolkit [24]. GEANT4 allows for full Monte-Carlo simulation of particle creation and propagation through any given medium. Using GEANT4 we propagate the core of the in-air particle shower, i.e. all particles within $1$ m of the shower axis, into an ice volume consisting of multiple horizontal 1-cm-thick layers of pure ice. Figure 16 shows the energy contained within a given radius from the shower core for two simulated proton-induced air showers at an altitude of $2.835$ km, which corresponds to a vertical atmospheric depth of $729$ g/cm3. Both showers were simulated using a zenith angle $\theta=0$ but differ by one decade in primary energy. For the shower with a primary energy of $E_{p}=10^{17}$ eV we see that at the given altitude the full air shower contains roughly $50$% of the primary energy. Within a radius of $1$ m from the shower core we find about $20$% of the primary energy, which is about half of the energy of the full particle shower at this altitude. A similar picture holds for the shower with a primary energy of $E_{p}=10^{18}$ eV, where the full shower contains close to $60$% of the primary energy and again half of that energy is contained within $1$ m of the shower axis. As illustrated by Figure 17 in Appendix A, propagating the showers through ice at the given altitude leads to high energy densities in the ice close to the shower core, dropping rapidly as the distance to the shower axis increases. Similar results were shown in Figures 5 and 6 in [22]. Furthermore, Figure 16 in [22] shows that we do not expect any significant changes in the total fluence of the in-ice radio emission around the Cherenkov cone, when the radius of the particle footprint being propagated into the ice is increased beyond $\sim 10$ cm. Important to note is that only for the in-ice cascade the simulation is limited to a radius of $1$ m from the shower axis. For the simulation of the in-air particle cascade and its corresponding radio emission, the whole air shower is accounted for.
Each layer of the ice has a constant density determined by the depth of the layer, following a typical polar ice density profile given by
@@
with $\rho_{\text{ice}}=917$ kg m-3, $\rho_{\text{surface}}=359$ kg m-3 and $t_{\text{firn}}=100$ m [41]. The ice density profile can be freely adjusted by the user. The radio emission is simulated using the endpoint formalism, using the implementation originally developed for the T-510 experiment at SLAC [42], with the addition of full ray tracing to account for the changing index of refraction in the ice. The index of refraction profile is not directly related to the layered density profile described by Equation 1, but instead follows a continuous exponential profile given below. Note that this means that the layered density profile of the ice volume only influences the development of the particle cascade, and does not influence the radio propagation. The features of the density profile at depths beyond $\sim 20$ m are therefore not relevant. The GEANT4-based particle propagation program has been discussed in detail in [22, 36].
As shown in [25], the endpoint formalism naturally includes the emission of transition radiation when the calculations of the emission in the first medium and the emission in the second medium are performed separately, which is how it is implemented in FAERIE.
@@ III Ray tracing
For the results presented in this work, the density of the atmosphere is described by a five-layer model for the South Pole fitted to a database used by the National Center for Environmental Predictions Global Forecast System in weather forecasting. Based on this fit, a five-layer model of the refractive index is created. Both the five-layer density model as well as the corresponding five-layer refractive index model are then used for the CORSIKA/CoREAS simulation [39]. The details of the models are given in Appendix B. In the following, the refractive index model for ice is set to
@@
corresponding to the model used by the Askaryan Radio Array at the South Pole outlined in [43]. More detailed information on the analytical ray tracer can be found in [44, 45].
Air-to-ice ray tracing describes ray bending in air, followed by refraction at the air-ice boundary and further bending in the ice. The ray-bending effect is less pronounced in air than in ice since the air refractive index value remains close to unity throughout. The refractive index profile is however still relevant, as the refractivity $R=n-1$ starts around $R=0$ at the top of the atmosphere and increases to approximately $R=3\times 10^{-4}$ at sea level, which can lead to a so-called refractive displacement for inclined air showers [46]. The polar ice refractive index changes from $n=1.35$ at the surface to $n=1.78$ within the first $100-200~{}$m. Therefore, significant refraction occurs near the ice surface due to a sharp change in the refractive index values. For ice-to-ice ray propagation, in general two solutions are found for each emitter-receiver pair. The first solution we will call the direct ray, which out of the two solutions has the shortest ray path between emitter and receiver. The second solution we will call the indirect ray, which can either be a refracted ray with a ray path longer than that of the direct ray, or a ray reflecting on the ice-air boundary. All three types of rays are shown in Figure 1.
@@ III.1 Interpolation of ray tracing parameters
A typical execution of the analytic ray tracing functions involving air and ice takes around $0.05-0.1$ ms. This is relatively fast but not fast enough for simulating a complete cosmic ray shower. A typical cosmic-ray shower will consist of $\mathcal{O}(10^{9})$ particles at EeV energies. In this case, propagating rays from each shower particle at each step to each antenna could take weeks or months, depending on the shower energy and the number of in-ice antennas. Therefore, it is not feasible to use the analytic ray tracing functions directly, and we have to move toward interpolation tables. Interpolation of ray tracing parameters from premade tables makes the ray tracing process significantly faster.
@@
To reduce memory usage during the simulation, antennas at the same depth in the ice share a single interpolation table for the in-air ray tracing. Since the index of refraction profile only depends on depth, the ray tracing applied is cylindrically symmetric. A grid covering a sufficiently large interval of horizontal distances can therefore be used by multiple antennas located at the same depth in the ice. For the in-air ray tracing the grid depth is varied from the ice layer altitude, generally around 3 km, up to 100100100~{}100km. The horizontal distance range for the grid is set by varying the ray launch angle $\theta$ from $89.9^{\circ}$ (almost horizontal) to $0^{\circ}$ (vertically down). Small variations in $\theta$ can cause almost exponential changes in the horizontal distance.
The in-ice ray tracing interpolation tables span intervals of 20 m on both axes, which covers the typical dimension of in-ice cascade. Since in ice the refractive index changes much more than it does in air, the in-ice grids are denser than the in-air grids. For in-ice ray tracing, one interpolation table is made for every single antenna, irrespective of the antenna depths in the ice. A visualization of an interpolation table used for the in-ice ray tracing is given in Figure 2.
@@
Figure 2: A visualization of an interpolation table used for the in-ice ray tracing, for a given receiver position. Interpolation tables for the in-air ray tracing follow a similar structure, but instead cover a much larger area in air. A single in-air table is used for multiple receivers at the same depth in the ice.
@@ III.2 Modification of the endpoint formalism
Left: Visualization of several rays traced from the receiver to a line of emitters, including an air-ice boundary at $z=0$. Direct rays are shown in blue. Reflected rays are only considered when total internal reflection occurs and are shown in red. Transmitted rays are shown in green. Right: the boost factor estimations along with the numerically computed value of $\frac{dt}{dt^{\prime}}$ as a function of depth. The diamonds correspond to the calculations using $\hat{r}$ pointing directly from emission point to receiver point. The stars correspond to the values of the boost factor when evaluating $n$ at the emission point and using the launch direction for $\hat{r}$. The dots correspond to the numerically calculated values of the boost factor. The different regions are associated with direct (blue), reflected (red) and transmitted rays (green).
The endpoint formalism calculates the electric field observed at a position $\vec{x}$ emitted by a charged particle taking a propagation step (track segment) in the simulation by using the formula
@@
where the plus sign is applied using the start point of the step and the minus sign using the end point of the step [25]. The combination of both then gives the net contribution of the step. Here, $\Delta t$ is the sampling time interval of the observer, $q$ the charge of the particle, $\hat{r}$ the direction from the start/end point toward the antenna, $\vec{\beta}^{*}$ the velocity of the charge during the step, $R$ the distance between the start/end point and the antenna, and $n$ the index of refraction of the medium at the respective emission point. The variable $t$ refers to the arrival time of the emission at the observer, and relates to the emission time $t^{\prime}$ via $t=t^{\prime}+nR/c$. The emission time $t^{\prime}$ is often referred to as the retarded time.
The so-called boost factor $B=1-n\vec{\beta}^{*}\cdot\hat{r}$ found in Equation 3 arises from classical theory and can more generally be described as $B=\frac{dt}{dt^{\prime}}$. The implementation as used in the endpoint formalism considers signals traveling along straight lines, i.e. $\hat{r}$ points directly from emission point to receiver point. This approach is however not valid when working in media with a nonconstant index of refraction where formalisms such as ray tracing need to be used to correctly describe signal propagation.
Using a ray tracer calculating the derivative $\frac{dt}{dt^{\prime}}$ numerically, we found that the relation $B=1-n\vec{\beta}^{*}\cdot\hat{r}$ still holds provided that $n$ is evaluated at the emission point and $\hat{r}$ is interpreted as the launching direction of the ray [47]. Figure 3 shows a visualization of the ray tracing together with the comparison of different estimators of the boost factor $B$ to the numerically computed value of $\frac{dt}{dt^{\prime}}$. Rays were traced from a receiver in the ice to a line of emitting points, including an air-ice boundary at $z=0$. Both the in-air and the in-ice medium have an associated exponential index of refraction profile. Evaluating the boost factor using a straight line calculation clearly diverges from the numerically calculated value, while evaluating $n$ at the emission point and using the launch direction for $\hat{r}$ agrees well for all direct, reflected and refracted rays.
@@
Finally, a rotation of the electric field is applied in the plane of the ray. This rotates the electric field such that its component in the plane of the ray is perpendicular to the receiver direction, instead of the launching direction. A visualization of the endpoint formalism including ray tracing is shown in Figure 4.
Figure 4: A visualization of the variables used in the endpoint formula including ray tracing (Equation 3), in the specific case where the start point, end point and receiver coincide in the same plane. The circles represent the start and end point of the emitter step, the square represents the receiver. All vectors shown in the illustration lie in the same plane.